# Surface Tension of a Topological Phase

### v2

@@ Abstract
Growth of clusters at finite temperature: At finite $T$, thermal fluctuation reduces the stiffness of domain wall thereby reducing $\sigma_{\rm{cl}}$ from its $T=0$ value ($2J$[27]. It is useful to note we work with $T<T_{c}$ where classical transition temperature $T_{c}\approx 2.269J$ [55] (where Boltzmann constant is set to unity). Moreover, we consider temperatures much smaller than fermionic energy-scales, thus the effect of thermal fluctuation on quantum surface tension can be ignored. Thus with increasing temperature, since effective $\sigma_{\rm{cl}}$ reduces, we expect that $\gamma$ increases [61].
For numerical computation of $n_{c}^{\rm{cl}}$ and $n_{c}^{\rm{Q}}$ at finite $\beta=1/T$, we resort to Monte Carlo (MC) simulation where both the spins and fermions contribute to total energy of any configuration. However, given electronic time-scales are much faster than the spin time-scales, the electronic effects can be effectively included as internal energy corrections to the spin configurations within Metropolis algorithm [62] (see SM [61]). We start from an initial configuration with $n=l^{2}$ up-spins while rest of the spins are down. We then investigate the quantity
@@
where $P_{+}(n)$ and $P_{-}(n)$ count the number of MC steps in which size $n$ of a cluster increases to $(n+1)$ and decreases to $(n-1)$ respectively [6]. Critical cluster-sizes $n_{c}^{\rm{cl}}$ for classical situation ($\kappa=0$) and $n_{c}^{\rm{Q}}$ for both topological ($\kappa=1.2$) and trivial ($\kappa=2.5$) situations are then obtained from when $\nu$ crosses 00 [63], where we find $n_{c}^{\rm{Q}}>n_{c}^{\rm{cl}}$ in topological situation and $n_{c}^{\rm{Q}}\approx n_{c}^{\rm{cl}}$ in trivial situation (see Fig. 3(a)) thus showing the characteristic rise of surface tension only in the topological case.
To observe the growth (shrinkage) of clusters with time $t$, we calculate average local magnetization $\langle s_{x,y}\rangle$ for any unit cell at position ${\textbf{r}}=(x,y)$ as a function of $t$. The behavior of $\langle s_{x,y}\rangle$ for different values of $\kappa$ and $t$ is shown in Fig. 3(b). Our study reveals that when initial size of cluster (having up-spins) $n>n_{c}^{\rm{Q}}$, the cluster grows with time and when $n_{c}^{\rm{cl}}>n$, the cluster shrinks. This behavior remains unchanged in both the classical ($\kappa=0$) and the topological ($\kappa=1.2$) situations. However, when $n_{c}^{\rm{Q}}>n>n_{c}^{\rm{cl}}$, although the cluster grows for $\kappa=0$, it shrinks for $\kappa=1.2$ (see Fig. 3(b)), thus establishing the role of topological surface tension in altering growth (shrinkage) of clusters. Moreover, modifying the quench protocols can further allow tunability in such growth processes (see SM [61]).
@@
Figure 4: Edge-state evolution: Local density of states (LDOS) of edge-modes for initial cluster-sizes (a) $n=441$ ($n>n_{c}^{\rm{Q}}$), (b) $n=225$ ($n<n_{c}^{\rm{Q}}$) at times $t=0,40,80$ (in the units of Monte Carlo steps per site), where $n_{c}^{\rm{Q}}=304$. The parameters chosen here are: $J=0.2$, $T=1/\beta=0.33$ (where $T_{c}\approx 0.45$), $h=0.02$, $\kappa=1.2$. The number of unit cells in square lattice is $L^{2}=4096$. We observe that LDOS of edge-modes captures the signatures of growth (when $n>n_{c}^{\rm{Q}}$) and shrinkage (when $n<n_{c}^{\rm{Q}}$) of clusters for $20<|\kappa|<$ 2.
@@
Edge-state evolution: In order to observe the edge-states in a nucleating droplet, we take a typical spin configuration and solve the fermionic problem at every step in the MC evolution  [61] to obtain the local density of states (LDOS) defined as ${\rm{LDOS}}=\sum_{j,\alpha}|\langle i\alpha|\psi_{j}\rangle|^{2}$ (where $j$ runs over all eigenstates corresponding to single-particle energy-eigenvalues $|\epsilon_{j}|<0.05W$ near the Fermi energy where $W$ is the bandwidth of single-particle spectrum) at $i$-th unit cell with coordinate $(x,y)$ and orbitals $\alpha=A,B$ at various times $t$. Unsurprisingly the edge-state evolution in topological situation ($20<|\kappa|<$ 2) follows the signatures of growth and shrinkage of clusters. We find that edge-modes are indeed localized at the boundary of clusters, however, unlike the configuration averaged circularly symmetric clusters (see Fig. 3(b)) any typical configuration has a rugged domain wall between two different topological regions (see Fig. 4). When $n>n_{c}^{\rm{Q}}$, the boundary of cluster grows with time $t$, while it shrinks and eventually vanishes when $n<n_{c}^{\rm{Q}}$ (see Fig. 4(a,b)). Taken together, all these results show a characteristic surface tension which can be alluded to a topological phase.
@@
Outlook: In this work, we investigate nucleation of a scalar field when coupled to a topological fermionic field. Taking a concrete example of an Ising model coupled to a Chern insulator, we find that fermionic quantum contribution leads to an additional correction to surface tension of a nucleating cluster. Interestingly, the coupling parameter of the two fields can serve as a new tunable parameter for controlling the critical cluster-size in nucleation, thus affecting its growth and shrinkage. Ranging from implications on fundamental aspects of coupled topological field theories, just nucleation physics may also play a role in formation of topological domains such as Chern mosaic structures seen in strongly correlated topological systems [36]. Our work opens up a range of questions regarding physics of nucleation when coupled to higher dimensional topological systems where one can systematically pose the role of such surface physics in modifying critical properties of the classical fields. The imminent breakdown of adiabaticity would result in interplay of quantum and thermal fluctuations leading to intriguing critical theories with characteristic exponents. Furthermore, the role of Goldstone fluctuations when the classical field has continuous symmetry can also be interesting. The study of spinodal decomposition and coarsening dynamics [64, 65] could be important future directions.
@@
Acknowledgements: We acknowledge fruitful discussions with Amit Agarwal, Vijay B. Shenoy, Diptarka Das, Sabyasachi Chakraborty, Soumya Sur, Subrata Pachhal, Rahul Singh. S.M. acknowledges support from PMRF Fellowship, India. AA acknowledges support from IITK Initiation Grant (IITK/PHY/2022010). Numerical calculations were performed on the workstations Wigner and Syahi at IITK.
Supplemental Material to “Surface Tension of a Topological Phase”
I Sources of Surface Tension
@@
In this section we demarcate both the bulk and edge contributions to the surface tension $\Delta\sigma$ for the model discussed in the main text.
I.1 Calculation of edge mode energy (Analytical)
@@
$H=-M\sum_{i}{\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{}{\bf{\Psi}}_{i}-\sum_{\langle ij\rangle}({\bf{\Psi}}_{i}^{\dagger}{\large{\eta}}_{ij}~{}{\bf{\Psi}}_{j}+{\rm{h.c.}}).$ (S1)
For periodic boundary conditions in both $x$ and $y$ directions, we resort to momentum space where Hamiltonian for $k_{x},k_{y}\in[-\pi,\pi)$ is
$H_{k_{x},k_{y}}=\sin(k_{x})\sigma_{x}+\sin(k_{y})\sigma_{y}+(-M-\cos(k_{x})-\cos(k_{y}))\sigma_{z}.$ (S2)
@@
$E(k_{x},k_{y})=\pm\sqrt{\sin^{2}(k_{x})+\sin^{2}(k_{y})+(M+\cos(k_{x})+\cos(k_{y}))^{2}}.$ (S3)
Now, in a ribbon geometry with periodic (open) boundary condition in $x$ ($y$) direction, $k_{x}$ remains a good quantum number and hosts $x$ dispersing edge states on the top and bottom boundaries in the topological phase ($20<|M|<$ 2). The dispersion in the two situations (open and periodic boundary conditions) and the local density of states (LDOS) of the edge-states are shown in Fig. S1(a-c).
Figure S1: Spectrum and edge modes in BHZ model: (a) Single-particle energy spectrum as a function of $k_{x}$ for BHZ model with $M=1.0$ for periodic boundary in $x$ direction and open boundary in $y$ direction, (b) single-particle energy spectrum as a function of $k_{x}$ for BHZ model with $M=1.0$ for periodic boundary in both $x$ and $y$ directions, (c) Local density of states (LDOS) of edge-modes at unit cell with coordinate $(x,y)$ for BHZ model with periodic boundary in $x$ direction and open boundary in $y$ direction for $L=32$, $M=1.0$.
The difference in the ground-state energy for a half-filled system between the two situations (a) and (b) determines the mean edge-mode energy when the system is deep in the topological phase ($0\ll|M|\ll 2$). For instance, at $M=1$ an edge state wavefunction that decays exponentially in the $y$ direction has an effective energy $E_{\rm{edge}}(k_{x})\sim-\sin(k_{x})$ near $k_{x}=\pi$ (see Fig. S1(a)). However, such an edge state can only be defined appropriately until it mixes with the bulk modes. The valence bulk band maxima $E_{\rm{valence}}(k_{x})$ is determined by $k_{y}=\pi$ mode when $20<M<$ 2 (similarly by $k_{y}=0$ mode when $-2<M<0$). Thus the range of $k_{x}$ until which the edge states survive is given by $k^{c}_{x}=\cos^{-1}(-M+1)=\pm\frac{\pi}{2}$ for $M=1$. The edge state energy is thus given by the difference
[…]
For a $L\times L$ square lattice with periodic boundary condition in $x$ direction and open boundary condition in $y$ direction, the edge length $=2L$ and thus the edge-mode energy per unit length is
[…]
which can be evaluated exactly and is $=\frac{(\pi-2)}{4\pi}$ for $M=1$. For a general linearly dispersing edge state where the dispersion is given by
$E_{\text{edge}}(k_{x})\sim\hbar v_{F}k_{x}.$ (S6)
@@
[…]
In the present Hamiltonian, since $v_{F}\sim 1$, and thus the surface tension arising from the edge is expected to vary as the square of the bulk gap. In the next section, we compare this analytic estimate with the numerical results of edge and bulk contribution to the surface tension.
Figure S2: Contribution of edge-modes and bulk states to surface tension: (a) Surface tension $\Delta\sigma$, contributions of bulk states $\Delta\sigma_{\rm{bulk}}$ and edge-modes $\Delta\sigma_{\rm{edge}}$ to surface tension as a function of $\kappa$ where $\Delta\sigma_{\rm{bulk}}$ and $\Delta\sigma_{\rm{edge}}$ are obtained using Eq. (S11a) and Eq. (S11b) respectively. (b) Difference of energy $\Delta E_{j}$ of $j$-th single-particle level (see Eq. (S8)) below the Fermi energy in topological situation ($\kappa=1.2$) when the size of the droplet is $n=l^{2}$. (c) $\Delta E_{j}$ as a function of $j$ in trivial situation ($\kappa=2.5$) when $l=16$. (d) $\Delta\sigma_{\rm{edge}}$ and analytical estimate of $\Delta{\mathcal{E}}_{\rm{edge}}$ (see Eq. (S7)) as a function of $\kappa\equiv M$. In all plots, number of unit cells in square lattice is $L^{2}=4096$.
I.2 Contribution of bulk and edge states (Numerical)
In order to isolate the contributions to the surface tension due to the bulk and edge states separately, we devise the following numerical scheme. We compare the single particle energy eigenvalues between the two cases (i) a uniform square region of size $L^{2}$ with all $s_{i}=-1$ (i.e. before quench), (ii) the case where within the region of size $L^{2}$, a square-shaped droplet of size $n=l^{2}$ with $s_{i}=1$ is now formed (i.e. after quench). Each of these systems has $2L^{2}$ single particle energy eigenvalues which can be enumerated and indexed in ascending order from $j=1,\ldots,2L^{2}$. The difference in the single particle eigenvalues between the two cases is
$\Delta E_{j}=E_{j}(\text{after quench})-E_{j}(\text{before quench})$ (S8)
and their additive sum until the Fermi energy ($j=L^{2}$) is the essential contribution to surface tension $\Delta\sigma$ at temperature $T=0$ which was evaluated in the main text using
$\Delta E=\sum_{j=1}^{L^{2}}\Delta E_{j}=4\Delta\sigma\sqrt{n}.$ (S9)

### v3

@@ IV Growth of clusters at finite temperature
At finite $T$, thermal fluctuation reduces the stiffness of domain wall thereby reducing $\sigma_{\rm{cl}}$ from its $T=0$ value ($\sim 2J$[54, 17]. It is useful to note we work with $T<T_{c}$ where classical transition temperature $T_{c}\approx 2.269J$ [46] (where Boltzmann constant is set to unity). Moreover, we consider temperatures much smaller than fermionic energy-scales, thus the effect of thermal fluctuation on $\Delta\sigma$ can be ignored. Thus with increasing temperature, since effective $\sigma_{\rm{cl}}$ reduces, we expect that $\gamma$ increases (see Appendix C).
To analyze the critical cluster size and growth of the clusters at finite $T$, we resort to Monte Carlo (MC) simulation and employ the Metropolis algorithm [45]. When the spins and fermions are decoupled ($\kappa=0$), a random spin $s_{i}$ is chosen and the energy $\Delta E_{i}$ required to flip the spin (i.e., $s_{i}\to-s_{i}$) is calculated as:
$\Delta E_{i}=2Js_{i}\sum_{j\in{\mathcal{N}}_{i}}s_{j}+2hs_{i},$ (9)
where $j$ runs over all the nearest neighbors (${\mathcal{N}}_{i}$) of $i$-th spin. Now, if $\Delta E_{i}<0$ or if $p_{i}\leq\exp(-\beta\Delta E_{i})$ for a randomly chosen $p_{i}\in[0,1]$ where $\beta=1/T$, the spin flip is accepted and the $i$-th spin becomes $s_{i}$. If none of these conditions is satisfied, the spin-flip is rejected and $i$-th spin remains $s_{i}$.
When the classical Ising field is coupled to the fermionic field ($\kappa\neq 0$), both the spins and fermions contribute to total energy. Here we solve the fermionic problem to compute $\Delta E_{i}$ at each step of Monte Carlo simulation. In this method, we propose a spin flip $s_{i}\to-s_{i}$ at a randomly chosen $i$-th unit cell at each Monte Carlo step. We then explicitly calculate the difference $\Delta E_{i}^{\rm{Q}}$ in the ground state energies between the configurations after and before the proposed $i$-th spin-flip for the fermionic problem as:
$\Delta E_{i}^{\rm{Q}}=E^{\rm{Q}}\left(s_{1},...,-s_{i},...,s_{L^{2}}\right)-E^{\rm{Q}}\left(s_{1},...,s_{i},...,s_{L^{2}}\right),$ (10)
$\Delta E_{i}=2Js_{i}\sum_{j\in{\mathcal{N}}_{i}}s_{j}+2hs_{i}+\Delta E_{i}^{\rm{Q}}.$ (11)
Eq. (11) can be used to check the acceptance or rejection of the spin-flip at each Monte Carlo step when $\kappa\neq 0$. Under this protocol, the spin fluctuations and their effect on fermions are exactly accounted for.
Figure 5: Critical cluster-size and droplet growth: (a) $\nu$ as a function of cluster-size $n$ for classical ($\kappa=0$), topological ($\kappa=1.2$) and trivial ($\kappa=2.5$) situations (see Eq. (12)). $\nu=0$ denotes the critical sizes which are $n_{c}^{\rm{cl}}=145$, $n_{c}^{\rm{Q}}=304$, $n_{c}^{\rm{Q}}=158$ respectively. While in topological situation $n_{c}^{\rm{Q}}>n_{c}^{\rm{cl}}$, in trivial situation $n_{c}^{\rm{Q}}\approx n_{c}^{\rm{cl}}$. (b) Initial and final average local magnetization $\langle s_{x,y}\rangle$ when clusters of sizes, $n=441$ ($n>n_{c}^{\rm{Q}}$), $n=225$ ($n_{c}^{\rm{Q}}>n>n_{c}^{\rm{cl}}$), $n=81$ ($n_{c}^{\rm{cl}}>n$) are evolved. The final time $t=100$ is in units of MC steps per site. Upper (lower) panel shows $\kappa=0$ ($\kappa=1.2$) and $J=0.2$, $T=1/\beta=0.33<T_{c}\approx 0.45$, $h=0.02$, $L^{2}=4096$. Averaging is performed over $10^{3}$ simulations. For classical ($\kappa=0$) situation, the cluster grows when $n>n_{c}^{\rm{cl}}$ but in topological situation ($\kappa=1.2$), growth occurs only when $n>n_{c}^{\rm{Q}}$.
We perform an exact MC simulation for $\kappa\neq 0$, starting from an initial configuration with a cluster of up-spins and size $n=l^{2}$ within the region of all down-spins. For a typical realization, the local magnetization ($s_{x,y}$) for any unit cell at position ${\textbf{r}}=(x,y)$ as a function of time $t$ (in units of Monte Carlo steps per site or after $tL^{2}$ Monte Carlo steps where $L^{2}$ is the number of spins in square lattice) illustrates the growth of the cluster (see Fig. 3). The exact MC simulation, however, is numerically expensive, since at every MC step, a matrix diagonalization for the fermionic system needs to be performed. This step can be simplified by devising an effective Monte Carlo method where the electronic effects can be effectively included as internal energy corrections to the spin configurations within Metropolis algorithm (see Appendix D). This simplification follows from the fact that the electronic time-scales are much faster than the spin time-scales. As shown in Fig. 4, the effective MC simulation can capture the growth (shrinkage) of the clusters with various initial sizes $n$ for both the decoupled ($\kappa=0$) and topological ($\kappa=1.2$) situations. We also observe that the boundaries of the clusters follow very rugged trajectories during their growth or shrinkage.
In order to understand the behavior of the growth (shrinkage) of the clusters for the decoupled and topological situations, we proceed to the numerical evaluation of the critical cluster sizes $n_{c}^{\rm{cl}}$ and $n_{c}^{\rm{Q}}$ using the effective MC simulation. Here, a cluster of up-spins and finite size $n$ is placed within the metastable background of down-spins and is subjected to a non-equilibrium situation (due to a sudden quench of $h$). Considering various realizations of the simulation, we then investigate the quantity
@@
where $P_{+}(n)$ and $P_{-}(n)$ count the number of MC steps in which size $n$ of a cluster increases to $(n+1)$ and decreases to (n1)(n-1) respectively [40]. When the cluster-size $n$ is smaller than the critical cluster-size, the tendency of shrinkage of the cluster is greater than that of growth, implying $P_{+}<P_{-}$ and $\nu<0$, while when $n$ is greater than the critical cluster-size, the tendency of growth of the cluster is greater than that of shrinkage, implying $P_{+}>P_{-}$ and $\nu>0$. Therefore, when $n$ is equal to the critical cluster-size, $\nu=0$. Critical cluster-sizes $n_{c}^{\rm{cl}}$ for classical situation ($\kappa=0$) and $n_{c}^{\rm{Q}}$ for both topological ($\kappa=1.2$) and trivial ($\kappa=2.5$) situations are thus obtained from when $\nu$ crosses 0, where we find $n_{c}^{\rm{Q}}>n_{c}^{\rm{cl}}$ in topological situation and $n_{c}^{\rm{Q}}\approx n_{c}^{\rm{cl}}$ in trivial situation (see Fig. 5(a)) thus showing the characteristic rise of the interfacial line energy only in the topological case.
We then move on to explore the universal feature of growth (shrinkage) of the clusters through the dynamics of the local magnetization averaged over various realizations. The behavior of the average local magnetization $\langle s_{x,y}\rangle$ for both $\kappa=0$ and $\kappa=1.2$ is shown in Fig. 5(b). Unlike the rugged boundaries of the clusters in a typical realization, the averaging of the local magnetization over various realizations renders the clusters to appear circularly symmetric. We find that when initial size of cluster (having up-spins) $n>n_{c}^{\rm{Q}}$, the cluster grows with time $t$ and when $n_{c}^{\rm{cl}}>n$, the cluster shrinks. This behavior remains unchanged in both the classical ($\kappa=0$) and the topological ($\kappa=1.2$) situations. However, when $n_{c}^{\rm{Q}}>n>n_{c}^{\rm{cl}}$, although the cluster grows for $\kappa=0$, it shrinks for $\kappa=1.2$ (see Fig. 5(b)). These observations establish the role of the topological interface energy contribution in altering growth (shrinkage) of clusters.
While we have established the role of topological interface energy contribution in affecting the nucleation physics, modifying the quench protocols can allow further tunability in such growth processes (see Appendix E).
V Edge-state evolution
@@
Figure 6: Edge-state evolution: Local density of states (LDOS) of edge-modes for initial cluster-sizes (a) $n=441$ ($n>n_{c}^{\rm{Q}}$), (b) $n=225$ ($n<n_{c}^{\rm{Q}}$) at times $t=0,40,80$ (in the units of Monte Carlo steps per site), where $n_{c}^{\rm{Q}}=304$. The parameters chosen here are: $J=0.2$, $T=1/\beta=0.33$ (where $T_{c}\approx 0.45$), $h=0.02$, $\kappa=1.2$. The number of unit cells in square lattice is $L^{2}=4096$. We observe that LDOS of edge-modes captures the signatures of growth (when $n>n_{c}^{\rm{Q}}$) and shrinkage (when $n<n_{c}^{\rm{Q}}$) of clusters for $0<|\kappa|<2$.
In order to observe the edge-states in a nucleating droplet, we take a typical spin configuration and solve the fermionic problem at every step in the MC evolution to obtain the local density of states (LDOS) defined as
${\rm{LDOS}}=\sum_{j}\left(|\langle i,A|\psi_{j}\rangle|^{2}+|\langle i,B|\psi_{j}\rangle|^{2}\right),$ (13)
@@
at $i$-th unit cell with coordinate $(x,y)$ (where $j$ runs over all eigenstates corresponding to single-particle energy-eigenvalues $|E_{j}|<0.05W$ near the Fermi energy where $W$ is the bandwidth of single-particle spectrum) at various times $t$. Unsurprisingly the edge-state evolution in topological situation ($0<|\kappa|<2$) follows the signatures of growth and shrinkage of clusters. We find that edge-modes are indeed localized at the boundary between two different topological regions in any typical realization (see Fig. 6). When $n>n_{c}^{\rm{Q}}$, the boundary of cluster grows with time $t$, while it shrinks and eventually vanishes when $n<n_{c}^{\rm{Q}}$ (see Fig. 6(a,b)). Taken together, all these results show a characteristic interfacial line energy which can be alluded to a topological phase.
VI Outlook
@@
In this work, we investigate nucleation of a scalar field when coupled to a topological fermionic field. Taking a concrete example of a two-dimensional Ising model coupled to a Chern insulator, we find that fermionic quantum contribution leads to an additional correction to the interfacial line energy in a nucleating cluster. Interestingly, the coupling parameter of the two fields can serve as a new tunable parameter for controlling the critical cluster-size in nucleation, thus affecting its growth and shrinkage. Ranging from implications on fundamental aspects of coupled topological field theories, just nucleation physics may also play a role in formation of topological domains such as Chern mosaic structures seen in strongly correlated topological systems [33]. Our work opens up a range of questions regarding physics of nucleation when coupled to higher dimensional topological systems where one can systematically pose the role of surface tension in modifying critical properties of the classical fields. The imminent breakdown of adiabaticity would result in interplay of quantum and thermal fluctuations leading to intriguing critical theories with characteristic exponents. Furthermore, the role of Goldstone fluctuations when the classical field has continuous symmetry can also be interesting. The study of spinodal decomposition and coarsening dynamics [18, 16] could be important future directions.
Acknowledgements.
@@
We acknowledge fruitful discussions with Amit Agarwal, Vijay B. Shenoy, Diptarka Das, Sabyasachi Chakraborty, Soumya Sur, Subrata Pachhal, Rahul Singh. S.M. acknowledges support from PMRF Fellowship, India. AA acknowledges support from IITK Initiation Grant (IITK/PHY/2022010). Numerical calculations were performed on the workstations Wigner and Syahi at IITK.
Appendix A Sources of Interface energy
@@
We demarcate both the bulk and edge contributions to $\Delta\sigma$ for the model discussed in the main text.
A.1 Calculation of edge mode energy
@@
$H=-M\sum_{i}{\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{\bf{\Psi}}_{i}-\sum_{\langle ij\rangle}({\bf{\Psi}}_{i}^{\dagger}{\large{\eta}}_{ij}~{\bf{\Psi}}_{j}+{\rm{h.c.}}).$ (14)
For periodic boundary conditions in both $x$ and $y$ directions, we resort to momentum space where Hamiltonian for $k_{x},k_{y}\in(-\pi,\pi]$ is
$H_{k_{x},k_{y}}=\sin(k_{x})\sigma_{x}+\sin(k_{y})\sigma_{y}+(-M-\cos(k_{x})-\cos(k_{y}))\sigma_{z}.$ (15)
@@
$E(k_{x},k_{y})=\pm\sqrt{\sin^{2}(k_{x})+\sin^{2}(k_{y})+(M+\cos(k_{x})+\cos(k_{y}))^{2}}.$}} (16)
Figure 7: Spectrum and edge modes in BHZ model: (a) Single-particle energy spectrum as a function of $k_{x}$ for BHZ model with $M=1.0$ for periodic boundary in $x$ direction and open boundary in $y$ direction, (b) single-particle energy spectrum as a function of $k_{x}$ for BHZ model with $M=1.0$ for periodic boundary in both $x$ and $y$ directions, (c) Local density of states (LDOS) of edge-modes at unit cell with coordinate $(x,y)$ for BHZ model with periodic boundary in $x$ direction and open boundary in $y$ direction for $L=32$, $M=1.0$.
Now, in a ribbon geometry with periodic (open) boundary condition in $x$ ($y$) direction, $k_{x}$ remains a good quantum number and hosts $x$ dispersing edge states on the top and bottom boundaries in the topological phase ($0<|M|<2$). The dispersion in the two situations (open and periodic boundary conditions) and the local density of states (LDOS) of the edge-states are shown in Fig. 7(a-c).
The difference in the ground-state energy for a half-filled system between the two situations (a) and (b) determines the mean edge-mode energy when the system is deep in the topological phase ($0\ll|M|\ll 2$). For instance, at $M=1$ an edge state wavefunction that decays exponentially in the $y$ direction has an effective energy $E_{\rm{edge}}(k_{x})\sim-\sin(k_{x})$ near $k_{x}=\pi$ (see Fig. 7(a)). However, such an edge state can only be defined appropriately until it mixes with the bulk modes. The valence bulk band maxima $E_{\rm{valence}}(k_{x})$ is determined by $k_{y}=\pi$ mode when $0<M<2$ (similarly by $k_{y}=0$ mode when 2<M<0-2<M<0). Thus the range of $k_{x}$ until which the edge states survive is given by $k^{c}_{x}=\cos^{-1}(-M+1)=\pm\frac{\pi}{2}$ for $M=1$. The edge state energy is thus given by the difference
$\Delta E_{\rm{edge}}^{\rm{analytical}}=\frac{L}{\pi}\int_{k^{c}_{x}}^{\pi}dk_{x}\left(E_{\rm{edge}}(k_{x})-E_{\rm{valence}}(k_{x})\right).$ (17)
For a $L\times L$ square lattice with periodic boundary condition in $x$ direction and open boundary condition in $y$ direction, the edge length =2L=2L and thus the edge-mode energy per unit length is
$\Delta{\mathcal{E}}_{\rm{edge}}$ $=\frac{\Delta E_{\rm{edge}}^{\rm{analytical}}}{2L}$
$=\frac{1}{2\pi}\int_{k^{c}_{x}}^{\pi}dk_{x}\left(E_{\rm{edge}}(k_{x})-E_{\rm{valence}}(k_{x})\right),$ (18)
which can be evaluated exactly and is $\frac{(\pi-2)}{4\pi}$ for $M=1$. For a general linearly dispersing edge state where the dispersion is given by
$E_{\text{edge}}(k_{x})\sim\hbar v_{F}k_{x}.$ (19)
@@
$\Delta{\mathcal{E}}_{\rm{edge}}\sim\frac{1}{2\pi}\int_{0}^{\frac{\Delta_{\rm{bg}}}{2\hbar v_{F}}}dk_{x}\hbar v_{F}k_{x}\sim\frac{\Delta_{\rm{bg}}^{2}}{16\pi\hbar v_{F}}.$ (20)
In the present Hamiltonian, since $v_{F}\sim 1$, and thus the interfacial line energy arising from the edge is expected to vary as the square of the bulk gap.
A.2 Contribution of bulk and edge states to interface energy
Figure 8: Contribution of edge-modes and bulk states to interfacial line energy: (a) Correction to the interfacial line energy $\Delta\sigma$, the contributions of bulk states $\Delta\sigma_{\rm{bulk}}$ and edge-modes $\Delta\sigma_{\rm{edge}}$ as a function of $\kappa$ where $\Delta\sigma_{\rm{bulk}}$ and $\Delta\sigma_{\rm{edge}}$ are obtained using Eq. (24a) and Eq. (24b) respectively. (b) Difference of energy $\Delta E_{j}$ of $j$-th single-particle level (see Eq. (21)) below the Fermi energy in topological situation ($\kappa=1.2$) when the size of the droplet is $n=l^{2}$. (c) $\Delta E_{j}$ as a function of $j$ in trivial situation ($\kappa=2.5$) when $l=16$. (d) $\Delta\sigma_{\rm{edge}}$ and analytical estimate of $\Delta{\mathcal{E}}_{\rm{edge}}$ (see Eq. (20)) as a function of $\kappa\equiv M$. In all plots, number of unit cells in square lattice is $L^{2}=4096$.
In order to isolate the contributions to the interfacial line energy due to the bulk and edge states separately, we devise the following numerical scheme. We compare the single particle energy eigenvalues between the two cases (i) a uniform square region of size $L^{2}$ with all $s_{i}=-1$ (i.e. before quench), (ii) the case where within the region of size $L^{2}$, a square-shaped droplet of size $n=l^{2}$ with $s_{i}=1$ is now formed (i.e. after quench). Each of these systems has $2L^{2}$ single particle energy eigenvalues which can be enumerated and indexed in ascending order from $j=1,\ldots,2L^{2}$. The difference in the single particle eigenvalues between the two cases is
$\Delta E_{j}=E_{j}(\text{after quench})-E_{j}(\text{before quench})$ (21)
and their additive sum until the Fermi energy ($j=L^{2}$) is the essential contribution to $\Delta\sigma$ at temperature $T=0$ which was evaluated in the main text using
$\Delta E=\sum_{j=1}^{L^{2}}\Delta E_{j}=4\Delta\sigma\sqrt{n}.$ (22)